Circular No. 6567

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) H. Matthews, Joint Astronomy Centre, Hilo, and Herzberg Institute of

Astrophysics; and D. Jewitt, University of Hawaii, report on observations of several isotopes in the gas coma of C/1995 O1 with the James Clerk Maxwell Telescope: “Preliminary measurements from Feb. 16 UT yield H^{12}CN(4–3)/H^{13}CN(4–3) = 90 ± 10 and HC^{14}N(4–3)/HC^{15}N(4–3) = 299 ± 30. Both ratios include a modest correction for the effects of optical depth in the HCN line. We believe that this is the first determination of ^{14}N/^{15}N in a comet. In addition, on Feb. 23 we determined C^{32}S(5–4)/C^{34}S(5–4) = 27 ± 3. The CS line was optically thin. Within the uncertainties, these isotope ratios are compatible with the terrestrial values ^{12}C/^{13}C = 89, 14N/^{15}N = 270, and^{32}S/^{34}S = 24. The 362.0465-GHz transition of DCN(5– 4) was also sought. A 3σ upper limit to the line ratio HCN(4–3)/DCN(5–4) = 300 was found on Feb. 15, indicating that deuterium is not highly enriched in C/1995 O1 relative to the interstellar value.”

GRB 970111 F. Frontera, Universitá di Ferrara; E. Costa and L. Piro, Istituto As-

trofisica Spaziale, Frascati; L. A. Antonelli, BeppoSAX, Rome; W. Voges and T. Boller, Max-Planck-Institut für Extraterrestrische Physik, Garching; and J. Greiner, Astrophysikalisches Institut Potsdam, communicate: “A 1300-s ROSAT HRI observation on Jan. 16 of the location of the GRB 970111 has revealed three x-ray sources in the field-of-view; an image is

available at http://www.aip.de:8080/~jcg/grb970111.html. Only one

of these three is inside the reduced error box (Hurley et al., IAUC 6545): RX J152845+1944.5, at α = 15^{h}28^{m}45^{s}.0, δ = +19^{o}44′31′′ (equinox 2000.0, error radius 10′′). This source coincides with the BeppoSAX source ‘a’ (IAUC 6539), the radio source VLA J1528.7+1945 (IAUC 6545), and the two galaxies (IAUC 6559); its intensity of 0.0082 ± 0.0025 count/s is consistent with that of the ROSAT all-sky-survey (RASS) source 1 (IAUC 6539), implying no x-ray variability between 1991 and 1997. If RX J152845+1944.5 = BeppoSAX ‘a’ is related to one of the two galaxies S1 or S2 (IAUC 6559), its x-ray luminosity is of the order of 5 × 10^{44} erg/s in the range 0.1–2.4 keV, which is typical for an active galactic nucleus. The BeppoSAX source ‘b’, as well as the two RASS sources 2 and 3 (IAUC 6539), are not detected in the HRI observation, with marginal evidence for an intensity drop by a factor of two for source 3 between 1991 and 1997.”

1997 February 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green